In each of the following problems, three sets are described. One of the sets is not the same as the other two. In each case, find the set that is not like the others. (you do not need to prove, a very brief explanation is sufficient)

(a) A = {a + b | a ∈ N, b ∈ N}; B = {a − b | a ∈ N, b ∈ N}; C = N.

(b) A = {x^2 | x ∈ N}; B = {x^2 | x ∈ Z}; C = {x^2 | x ∈ R}.

(c) A = {4z + 2 | z ∈ Z}; B = {4z − 2 | z ∈ Z}; C = {2z + 4 | z ∈ Z}.

(d) A = {m | m ∈ Z ∧ m ≡ 4 (mod 3)};

B = {m | m ∈ Z ∧ m ≡ −4 (mod 3)};

C = {m | m ∈ Z ∧ m ≡ 14 (mod 3)}.
Answer #1

a) Here the given sets ame, A = {atbl aftl, ben? B = { ara lang bent عنيا see thay in the set A, ale verb ors be positive verby AB Ersetih хси, e=&r InER {uf Herrey in the set A As, all neth, So, the mergers be A of full nember Similarly of all andAIB never Contain but C Contaning o. mez dy AS m = 4 (mod 3)} {mme meanm=-4 (med 3)} - Emmez 333 B m = 14 (mod this we see th